OpenStudy (anonymous):

Giving Medals to each question Answered A company sells each product they make for $72.00. They pay $2,100.00 in monthly costs, such as rent and bills, as well as $60.00 in manufacturing costs per unit. If the company sold 572 units last month, what was their total profit? A. $39,024.00 B. $6,136.80 C. $32,220.00 D. $4,764.00

undeadknight26 (undeadknight26):

easy: 72 x 572 = 41184 60 x 572 = 34320 41184 - 34320 = 6864 6864 - 2100 = 4764. its D.

undeadknight26 (undeadknight26):

:D

OpenStudy (anonymous):

O l l---D /\ Thumbs up!

OpenStudy (studygurl14):

572 x = 72 = $41,184 60 x 572 = 34,320 41,184 - 2,100 - 34,320 = $4,764 It is D, like @undeadknight26 said

OpenStudy (anonymous):

thank u u all for the explanation i was confused was multiplying but wasn't getting either of the answers